Circular No. 8230

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 U3 (NEAT) K. Lawrence, Jet Propulsion Laboratory, reports the discovery by the

NEAT project of a comet (discovery observation below). Observations by J. Young at Table Mountain on Oct. 23.2 UT show a 3′′ coma with a short, broad, fan-shaped tail ∼ 8′′ long spanning p.a. 255^{o}–285^{o}.

2003 UT  α_{2000}  δ_{2000}  Mag.
Oct. 22.28573  2^{h}16^{m}17.^{s}40  +22^{o}00′06′′.7  18.9

The available astrometry (including prediscovery LONEOS observations from Oct. 17), the following preliminary orbital elements, and an ephemeris appear on MPEC 2003-U54.

T = 2003 May 1.854 TT  ω = 355.^{o}717
e = 0.56512  Ω = 347.493 }2000.0
q = 2.38435 AU  i = 6.857
a = 5.48280 AU n^{o} = 0.076771 P = 12.8 years

SUPERNOVAE 2003iy AND 2003iz The discoveries of two apparent supernovae have been reported from

unfiltered CCD images: 2003iy in NGC 6143 by K. Itagaki (Teppo-cho, Yamagata, Japan, 0.6-m reflector; via S. Nakano, Sumoto, Japan) and 2003iz in UGC 638 by B. Beutler and W. Li (LOSS/KAIT; cf. IAUC 8229).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003iy Oct. 20.39 16^{h}21^{m}43.^{s}22 +55^{o}05′05′′.2 16.8 8′′ E, 4′′ S
2003iz Oct. 21.4  1 02 06.75 +26 56 21.3 17.7 18′′ E, 46′′ S

Additional magnitudes by Itagaki for SN 2003iy (which was also absent on his images of 2003 Oct. 6 and 9): 2002 June 28, [18.5; 2003 July 15, [18.5; Oct. 22.364 UT, 16.9. Additional KAIT magnitudes for SN 2003iz: Oct. 4.4, [19.0; 14.4, 18.3; 22.4, 17.7. (The offset for SN 2003iz above is corrected from that on CBET 50.)

COMET C/2003 T3 (TABUR) Total-visual-magnitude and coma-diameter estimates: Oct. 16.49 UT,

11.9, 0′.9 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 16.94, 12.0, 0′.5 (A. Amorim, Florianopolis, Brazil, 0.14-m reflector).

2003 October 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green